# Episodic Memory Verbalization using Hierarchical Representations of Life-Long Robot Experience

### v1

@@ I Introduction
Verbalizing their own experiences is an important ability robots should have to improve natural and intuitive human-robot interaction [1, 2, 3, 4]. It involves summarization of and question answering (QA) about a robot’s past actions, observations and interactions, such as the dialog shown on the right of Fig. 1. Building a representation of an agent’s Episodic Memory (EM) [5] is crucial to enable such verbalizations, as a system must efficiently store the information from the continuous stream of experience, organize it, and retrieve relevant past events from its EM in response to a user’s query. This is particularly challenging as the time horizon of the EM grows.
@@
To avoid training a system, which typically entails collecting large amounts of multimodal experience data, previous works [7, 8] use language-based representations of the past which can be obtained from pretrained multimodal models. Given such a language-based representation of an agent’s history of episodic events, a straightforward way to perform QA is to pass the question and the history to a large language model (LLM), and prompt it to produce an answer. While this works nicely for short histories [7, 9], in this paper, we focus on how to scale such approaches for verbalization of life-long experience streams. Although recent LLMs offer increasingly long context windows (i. e., the maximum number of tokens they can process), up to 2M tokens [10], previous studies [11, 12] have shown that these models have difficulty in using all information contained in such long contexts. Furthermore, the computation of transformer models scales quadratically with context length – reducing the number of tokens is thus time- and cost-effective.
@@
Therefore, to scale EMV to life-long experience streams while maintaining a low token budget, we propose to derive a tree-like representation from EM and use an LLM agent for QA to interactively search the tree to find relevant information. Our system, H-Emv (Hierarchical Episodic Memory Verbalization, Fig. 1), processes the continuous stream of experiences and inserts it into a hierarchical representation of the robot’s history of episodic events. Different levels of this hierarchy represent different abstraction levels, with the lowest level being raw observations and proprioception and higher levels being represented as natural language concepts. An LLM is prompted for segmentation and summarization in order to recursively create higher-level abstractions. To process queries to the EM, we repurpose the interactive prompting scheme described in our previous work [13]. An LLM is provided with the user’s query and functions to access the history tree, and the LLM itself decides which functions to use in order to fulfill the query, i. e., to answer the question or provide a summary. Since the history tree will grow large over time, we apply an interactive semantic search, inspired by work on robot navigation [14]. Specifically, we contract the tree, i. e., the LLM first sees only the top-level node, and then interactively explores the graph to retrieve the information relevant to the query.
To evaluate our system, we use simulated household episodes from TEACh [15], real-world egocentric human recordings from Ego4D [16], and real-world robot episodes from Armar-7, the newest member of the Armar humanoid robot family developed at KIT [17]. Our experiments show that H-Emv efficiently scales to extremely long histories of multiple simulated months or real-world human egocentric videos of over six hours, outperforming several baselines and ablations. Real-world robot demonstrations showcase the applicability of our system. We provide our code, evaluation data, and demonstration videos at hierarchical-emv.github.io.
@@ II Related Work
Robot Experience Verbalization:   The first work to introduce the term of “verbalizing” robot experiences was [2]. With a rule-based system, they converted a navigation route taken by a mobile service robot to natural language. [3] adapted this framework to verbalization of manipulation activities performed by a humanoid household robot. Similarly, [25] use templates to convert their robot’s observations and actions to natural language. More recent works phrase EMV in a more interactive setting, defined as summarization and QA on robot experiences [4]. Both [1, 6] propose end-to-end trained networks receiving multimodal experiences and a question to produce an answer. While [6] work on visual data only, [1] additionally use symbolic and subsymbolic information from the robot’s task execution and perception components. Both train on data from simulated household tasks. In contrast to these systems, H-Emv uses pretrained foundation models and does not require additional training data, thus increasing its versatility and easing deployment to the real world. Similar to our setting, QA from streaming data [26, 27] tackles the problem of answering questions based on a long stream of data, where the question is not known in advance and the raw data cannot be stored. However, we apply this to robotics, and approach it with an interpretable, modular system, instead of end-to-end trained memory models.
@@
Recent methods for VideoQA can be grouped into (i) end-to-end approaches [31, 32, 33, 9, 34] that typically connect pretrained frozen visual encoders with LLMs by some trained adapter, and (ii) training-free “socratic” [7] approaches [35, 36, 37, 38, 39, 40, 41] that invoke various off-the-shelf models to convert the video into text to be processed by a few-/zero-shot prompted LLM. For instance, [39, 41] use video captioning to produce a transcript of the video and then apply an LLM for QA based on this transcript. VideoTree [36] adaptively selects the frames to caption using a top-down query-relevance-based tree expansion instead of uniform sampling. [40] generate executable Python code from a question, invoking different APIs to query visual and language foundation models. MoReVQA [38] decomposes this into multiple stages, making the LLM’s job easier at each stage by focusing on either event parsing, grounding, or reasoning, instead of all at once. In contrast to these predefined prompting schemes, both [35, 37] use an LLM as an agent to analyze the video content in an interactive loop. While [37] iteratively ask the LLM whether to gather more detailed information (by captioning more intermediate frames) or produce the final answer, [35] provide the LLM with API functions invoking tools to search in a database of tracked objects or a memory of frame captions.
Our method similarly treats the LLM as an agent, thus not relying on any predefined information flow. However, we use the full flexibility of code [42] instead of single API calls like in [35, 37]. Compared to VideoTree [36], our history tree is constructed independently of the user’s query, since future questions cannot be known in advance in realistic settings, and storing lifelong “raw” video experiences is prohibitive [18]. In contrast to all of the above works, we consider real-world dates and times an integral part of the process. While the recent work TimeChat [43] is also time-sensitive, they refer to video timestamps instead of real-world date-times. Furthermore, and most crucially, we deal with long sequences of multimodal robotic experiences, with the longest experiment having over six hours of video or nearly two months on a simulated timeline.
@@ III-A Episodic Memory Construction
L1 – Scene Graphs: The first level of non-leaf nodes in the history tree has a one-to-one mapping to the L0 leafs. On this level, we derive a scene graph from the given observations, consisting of the detected objects as nodes and their spatial relations (e. g., on top, inside) as edges. The exact method for constructing the scene graph varies in our experiments. For the pure vision-based approach, objects are detected using pretrained models and heuristics are applied to infer semantically meaningful relations [8]. In our real-robot experiments, we use the existing components in our robot software framework ArmarX [44] that already provide semantic scene information.
@@
L4+ – Higher-Level Summaries: Summaries are generated dynamically by recursively asking an LLM to summarize the previous level’s nodes. Specifically, given the set of nodes $S_{\ell}$ at level $\ell\geq 3$, we list them in order and prompt an LLM to identify consecutive ranges of items that belong together considering their times and content, and provide a summary for each range. The output is parsed to group the child nodes and create the summary nodes $S_{\ell+1}$ for the next level. We apply this strategy recursively, until $|S_{\ell}|=1$ or there is no further reduction, i. e., $|S_{\ell+1}|=|S_{\ell}|$. In the latter, the LLM is explicitly prompted to provide a single concise summary of all items to force obtaining one root node.
@@ III-B Episodic Memory Access
Fig. 3 illustrates typical steps the LLM performs to answer a user’s question. Given the initially collapsed tree, the LLM first expands the root node’s children based on the requested date. It then selectively explores the respective child nodes that seem relevant to the question using the search function. Note that the LLM is prompted to collapse irrelevant nodes again in order to save token budget and speed up further requests. In the given example, when reaching a leaf node, the answer to the question is not evident from any of the natural-language summaries on each level, so the LLM decides to invoke a VLM to gather more information. Finally, it invokes the answer function to answer the user’s question.
Our implementation of the LLM agent uses a prompting style inspired by the simulated Python console of [13]. The LLM can issue any command – including compound statements such as loops – using the provided API. After the execution of the respective code, the LLM can “see” the output of its command(s), or any execution error. This process is repeated, and the prompt to the LLM always contains the (growing) execution history. Zero-shot experiments prompt the LLM with only a static prefix to explain the task and the available API, while few-shot prompting adds top-$k$ examples selected based on semantic similarity to the current user query. The final part of the prompt is always a string representation of the history tree’s current state.
@@ IV-A Simulated Household Episodes
[…]
@@
Gemini 1-pass FS all 26.5 50.3 29 27 325 19.0 34.8 16 27 663 11.6 32.6 17 28 824 5.3 21.4 8 23 1019 OOC
Gemini 1-pass FS 2nd 30.2 45.2 24 25 164 14.8 36.4 13 30 334 10.6 32.5 17 24 406 3.7 21.8 11 24 512 0.0 20.8 10 20 1256
H-Emv (1-shot) 2.4 18.9 12 24 7.2 2.0 17.9 14 24 13.6 1.9 16.8 16 27 14.2 2.4 17.5 16 24 10.5 0.1 17.6 16 18 16.4
@@
Gemini 1-pass FS 2nd 39.5 65.0 57 26 165 37.0 47.5 30 43 336 37.9 48.5 39 30 418 25.2 40.9 25 35 861 16.9 35.2 17 37 1591
H-Emv (1-shot) 4.3 32.9 27 38 11.6 3.9 28.7 36 37 10.5 2.2 30.3 36 35 11.8 5.8 26.3 29 35 10.6 1.0 17.3 24 33 15.1
@@
H-Emv (1-shot) 5.1 34.7 57 18 9.9 4.0 32.0 50 24 10.4 4.7 29.9 46 24 11.4 1.9 27.2 45 28 9.2 1.1 25.2 31 33 10.7
H-Emv (0-shot) 1.2 21.2 48 26 4.8 0.9 18.2 44 25 4.4 0.4 19.4 45 21 4.7 0.1 18.4 43 20 4.8 0.0 15.0 32 28 5.0
H-Emv (0-shot, L3) 0.9 19.9 24 29 25.2 1.0 17.6 18 3 65.6 0.1 16.1 25 24 60.4 0.1 12.1 17 24 82.9 0.0 4.4 4 11 16.6
Gemini 1-pass 0-shot L3 2.4 25.5 32 33 120 1.4 24.7 37 26 372 0.5 27.6 35 34 422 0.2 20.3 24 32 1055 0.0 15.2 3 5 1893
B: BLEU, R: ROUGE, $S_{c},S_{p}$: semantically categorized (partially) correct in %, T: number of 1K prompt token. Gray token costs exclude out-of-context (OOC) samples.
TABLE I: Results on Simulated Household Episodes from TEACh [15]
@@
Evaluation Metrics:   Evaluation of free-form EMV answers is hard since there can be many ways to formulate the correct answer, questions can be underspecified, and verifying abstract statements by grounding them in the history tree is a research question in itself. Following [1], we define a semantic categorization of a model’s hypothesis $h$ given the GT $g$ and question $q$: correct when $h$ is semantically equivalent to $q$; correctly summarized if $h$ is a correctly summarized version of $g$, still containing all relevant facts (in context of $q$); correct TMI (too much information) if $h$ is correct but overly specific; partially correct TMI if parts of $h$ are correct, but there are TMI parts and these are wrong; partially correct missing if parts of $h$ are correct, but relevant facts from $g$ are missing; wrong when $h$ could be an answer to $q$ but is none of the above; and no answer if $h$ is empty, completely irrelevant to $q$, or the model threw an error. Since categorizing each evaluated sample by hand is prohibitively expensive, we prompt GPT-4o [47] to perform this evaluation. For this, we started by annotating 60 samples by hand, and use these as a database to retrieve few-shot samples based on maximal marginal relevance [48]. We further tuned the prompts on a validation set of 100 hand-categorized model outputs. For reporting, we aggregate these semantic categories as the percentage of correct and partially correct samples, $S_{c}$ and $S_{p}$, respectively.
@@
We evaluate the agreement of the LLM’s predicted categories with manually annotated ones on 200 model results from our test data, resulting in an aggregated category accuracy of $88\%$, and per-class f-scores of $F_{1}(\text{correct})=0.89,F_{1}(\text{partially\_correct})=0.84,F_{1}(\text{wrong})=0.91$. The LLM categorizes correct and wrong samples very well and has the most difficulties on the partially correct labels. However, these categories are also defined imprecisely, and the inter-annotator agreement [49] between the first two authors has only a value of Cohen’s $\kappa=0.66$ ($n=110$), vs. $\kappa=0.91$ ($n=68$) when only considering correct/wrong. Thus, while not perfect, we use the LLM to automatically obtain reasonable score estimates.
In addition to manual and LLM-predicted categorization, we report standard automated metrics from machine translation that are often applied to free-form QA [6, 18]: BLEU-4 [50] and ROUGE-L (f-score) [51]. However, we emphasize that these surface-level metrics cannot grasp all varieties of correct answers in the EMV task, e. g., for a “when”-question, both of the following answers are correct, but have no word overlap at all: “at 4 PM” vs. “in the afternoon”.
@@
Settings and Baselines:   We evaluate under three settings: First, vision-only can act solely on the visual data stream. As a baseline, we prompt Gemini 1.5 Pro [10] in one pass with the sequence of images along with timestamps and the question. While we use GPT-4o for H-Emv, the 1-pass baseline uses Gemini because it requires extremely long requests. However, despite Gemini’s 2M token context length, we need to sample every 2nd frame for longer histories. We few-shot-prompt with one static example history of five episodes including ten QA samples for this history. H-Emv does not take raw images, but constructs history trees by inferring objects using YOLO-World [52] and actions using a LongT5 transformer model [53] fine-tuned on TEACh train. Second, vision + speech enriches the visual information with the dialog data from TEACh episodes, representing natural language commands given to the robot. This is simply added to the prompt for Gemini, and inserted into the history tree for H-Emv. Finally, full multimodal uses the recorded (GT) actions and goals from the TEACh episodes, as this information is typically available when a robot executes some actions. This setting also uses GT object information to compare system performance assuming perfect vision components. We compare H-Emv with one-shot and zero-shot prompting of the LLM agent. For preparing the few-shot samples, we use histories built from episodes in TEACh train, and record traces of manually using the Python console interface and the defined API to interact with the history tree until the given GT answer becomes evident. While we collect two to three samples per question type this way, making up 21 samples in total, we select only the top-1 sample when prompting the LLM, based on semantic similarity of the user’s questions. Semantic similarity is determined after asking gpt-4o-mini to cross out the task-specific words from the question so that an example from the same question type is retrieved (instead of an irrelevant example just mentioning the same objects or activities). We further ablate the hierarchical summarization: H-Emv 0-shot L3 is our method without LLM-generated summaries (L4+), still using the interactive agent to explore an initially collapsed list of L3 nodes. Last, Gemini 1-pass 0-shot L3 is a baseline presenting the fully expanded tree (L3 and lower) to the LLM along with the question in a single prompt, thus not using the LLM as an agent.
Results:   Results of our TEACh experiments can be found in Table I. First, we can observe that every method’s performance decreases with increasing $|h|$. Further, comparing the surface metric scores (B, R) of vision-only 1-pass with full multimodal H-Emv demonstrates that these metrics are not sufficient for evaluating EMV: While vision-only 1-pass has significantly higher B and R, H-Emv actually performs better on the semantic scores. This can be explained by our 1-pass prompting seeing more QA samples and thus better picking up the vocabulary of the generated data, which does not necessarily improve the correctness of the hypotheses but increases n-gram overlap. In contrast, H-Emv using GPT-4o tends to give more free-form answers, especially since it uses hierarchically generated summaries.
@@
Focusing on the vision-only and vision + speech results, the Gemini 1-pass baseline outperforms H-Emv for shorter histories. This is reasonable, as the baseline can directly access the full stream of visual information, whereas our hierarchical system suffers from error propagation and is limited by pretrained vision components. In particular, the history tree could contain incomplete or wrong information or our method could fail by expanding the wrong nodes of the tree, which cannot happen to the 1-pass baseline. However, token costs scale linearly with history length for 1-pass, while it stays approximately constant for H-Emv. The performance also drops faster for 1-pass, with H-Emv reaching comparable or better semantic scores for $|h|\geq 25$.
In contrast, when circumventing the limitations of perception components by using GT object detection and action information (full multimodal setting), H-Emv outperforms the 1-pass system in the semantic metrics, with a token budget two orders of magnitudes smaller. Further, 1-shot prompting significantly helps, but 0-shot also works reasonably well with half the token costs. Ablating the hierarchical higher-level summaries significantly increases token cost and leads to worse performance, also resulting in OOC errors for longer histories (when the LLM expands all nodes).

### v2

@@ I Introduction
Verbalizing their own experiences is an important ability robots should have to improve natural and intuitive human-robot interaction [1, 2, 3, 4, 5]. It involves summarization of and question answering (QA) about a robot’s past actions, observations and interactions, such as the dialog shown on the right of Fig. 1. Building a representation of an agent’s Episodic Memory (EM) [6] is crucial to enable such verbalizations, as a system must efficiently store the information from the continuous stream of experience, organize it, and retrieve relevant past events from its EM in response to a user’s query. This is particularly challenging as the time horizon of the EM grows.
@@
To avoid training a system, which typically entails collecting large amounts of multimodal experience data, previous works [8, 9] use language-based representations of the past which can be obtained from pretrained multimodal models. Given such a language-based representation of an agent’s history of episodic events, one practical way to perform QA is to pass the question and the history to a large language model (LLM) and prompt it to produce an answer. While this works nicely for short histories [8, 10], in this paper, we focus on how to scale such approaches for verbalization of life-long experience streams. Although recent LLMs offer increasingly long context windows (i. e., the maximum number of tokens they can process), up to 2M tokens [11], previous studies [12, 13] have shown that these models have difficulty in using all information contained in such long contexts. Furthermore, the computation of transformer models scales quadratically with context length – reducing the number of tokens is thus time- and cost-effective.
@@
Therefore, to scale EMV to life-long experience streams while maintaining a low token budget, we propose to derive a tree-like representation from EM and use an LLM agent for QA to interactively search the tree to find relevant information. Our system, H-Emv (Hierarchical Episodic Memory Verbalization, Fig. 1), processes the continuous stream of experiences and inserts it into a hierarchical representation of the robot’s history of episodic events. Different levels of this hierarchy represent different abstraction levels, with the lowest level being raw observations and proprioception and higher levels being represented as natural language concepts. An LLM is prompted for segmentation and summarization in order to recursively create higher-level abstractions. To process queries to the EM, we repurpose the interactive prompting scheme described in our previous work [14]. An LLM is provided with the user’s query and functions to access the history tree, and the LLM itself decides which functions to use in order to fulfill the query, i. e., to answer the question or provide a summary. Since the history tree will grow large over time, we apply an interactive semantic search, inspired by work on robot navigation [15, 16]. Specifically, we contract the tree, i. e., the LLM initially only sees the top-level node, and then interactively explores the graph to retrieve the information relevant to the query.
To evaluate our system, we use simulated household episodes from TEACh [17], real-world egocentric human recordings from Ego4D [18], and real-world robot episodes from Armar-7, the newest member of the Armar humanoid robot family developed at KIT [19]. Our experiments show that H-Emv efficiently scales to extremely long histories of multiple simulated months or real-world human egocentric videos of over six hours, outperforming several baselines and ablations. Real-world robot demonstrations showcase the wide applicability of our system. We provide our code, evaluation data, and demonstration videos at hierarchical-emv.github.io.
@@ II Related Work
Robot Experience Verbalization:   The first work to introduce the term of “verbalizing” robot experiences was [2]. With a rule-based system, they converted a navigation route taken by a mobile service robot to natural language. [3] adapted this framework to verbalization of manipulation activities performed by a humanoid household robot. Similarly, [27] use templates to convert their robot’s observations and actions to natural language. More recent works phrase EMV in a more interactive setting, defined as summarization and QA on robot experiences [4]. Both [1, 7] propose end-to-end trained networks receiving multimodal experiences and a question to produce an answer, using training data from simulated household tasks. While [7] work on visual data only, [1] additionally use symbolic and subsymbolic information from the robot’s task execution and perception components. In contrast to these systems, H-Emv uses pretrained foundation models and does not require additional training data, thus increasing its versatility and easing deployment to the real world. Most similar to our work, ReMEmbR [28] propose to use a robot’s experience for QA and navigation. Using an LLM as an agent to decide what data to retrieve from a vector store of memories, they can answer queries based on time, place, and content. The “search” function available to H-Emv’s LLM agent works similar to this but acts on a specific tree node instead of the overall history. Further, ReMEmbR’s memory is not hierarchical, thus limiting both temporal and content-based retrieval to short moments in time, struggling to handle queries that require larger context (such as “summarize your day”). Focusing on failure identification and recovery, [29] produce robot narrations from multimodal robotic data streams. This can be seen as orthogonal to our work as these narrations could enrich the lower levels of our hierarchy. Similar to our setting, QA from streaming data [30, 31] tackles the problem of answering questions based on a long stream of data where the question is not known in advance and the raw data cannot be stored. However, we apply this to robotics, and approach it with an interpretable, modular system instead of end-to-end trained memory models. Furthermore, some works in natural language processing propose hierarchical variants of retrieval-augmented generation to tackle multi-document QA tasks [32, 33, 34], which is similar to our work in that a graph-like structure is accessed, but differs as we use an LLM as an agent and apply it on inherently temporal, multimodal robotics data.
@@
Recent methods for VideoQA can be grouped into (i) end-to-end approaches [38, 39, 40, 10, 41] that typically connect pretrained frozen visual encoders with LLMs by some trained adapter, and (ii) training-free “socratic” [8] approaches [42, 43, 44, 45, 46, 47, 48] that invoke various off-the-shelf models to convert the video into text to be processed by a few-/zero-shot prompted LLM. For instance, [46, 48] use video captioning to produce a transcript of the video and then apply an LLM for QA based on this transcript. VideoTree [43] adaptively selects the frames to caption using a top-down query-relevance-based tree expansion instead of uniform sampling. Both [49, 50] build structured representations from egocentric video using various models, e. g., for object and hand tracking, as well as activity and location labeling. [47] generate executable Python code from a question, invoking different APIs to query visual and language foundation models. MoReVQA [45] decomposes this into multiple stages, making the LLM’s job easier at each stage by focusing on either event parsing, grounding, or reasoning, instead of all at once. In contrast to these predefined prompting schemes, both [42, 44] use an LLM as an agent to analyze the video content in an interactive loop. While [44] iteratively ask the LLM whether to gather more detailed information (by captioning more intermediate frames) or produce the final answer, [42] provide the LLM with API functions invoking tools to search in a database of tracked objects or a memory of frame captions.
Our method similarly treats the LLM as an agent, thus not relying on any predefined information flow. However, we use the full flexibility of code [51] instead of single API calls like in [42, 44]. Compared to VideoTree [43], our history tree is constructed independently of the user’s query, since future questions cannot be known in advance in realistic settings, and storing lifelong “raw” video experiences is prohibitive [20]. In contrast to all of the above works, we consider real-world dates and times an integral part of the process. While the recent work TimeChat [52] is also time-sensitive, they refer to video timestamps (“second 42 to 57”) instead of real-world date-times (“yesterday in the afternoon”). Furthermore, and most crucially, we deal with long sequences of multimodal robotic experiences, with the longest experiment having over six hours of video or nearly two months on a simulated timeline.
@@ III-A Episodic Memory Construction
L1 – Scene Graphs: The first level of non-leaf nodes in the history tree has a one-to-one mapping to the L0 leafs. On this level, we derive a scene graph from the given observations, consisting of the detected objects as nodes and their spatial relations (e. g., on top, inside) as edges. The exact method for constructing the scene graph varies in our experiments. For the pure vision-based approach, objects are detected using pretrained models and heuristics are applied to infer semantically meaningful relations [9], whereas in our real-robot experiments, we use the existing components in our robot’s software framework ArmarX [53] that already provide semantic scene information.
@@
L4+ – Higher-Level Summaries: Summaries are generated dynamically by recursively asking an LLM to summarize the previous level’s nodes. Specifically, given the set of nodes $S_{\ell}$ at level $\ell\geq 3$, we list them in order and prompt an LLM to identify consecutive ranges of items that belong together considering their times and content, and provide a summary for each range. The output is parsed to group the child nodes and create the summary nodes $S_{\ell+1}$ for the next level. We apply this strategy recursively, until $|S_{\ell}|=1$ or there is no further reduction, i. e., $|S_{\ell+1}|=|S_{\ell}|$. In the latter, the LLM is explicitly prompted to provide a single concise summary of all items to force obtaining one root node. To make this procedure robust to LLM errors, e. g., invalid output syntax, missing an item or including one in multiple ranges, we re-prompt on failure, listing detailed error messages and asking for improvement. In case the LLM fails to provide a valid grouping after three trials, we fall back to the single summary prompting.
@@ III-B Episodic Memory Access
Fig. 3 illustrates typical steps the LLM performs to answer a user’s question. Given the initially collapsed tree, the LLM first expands the root node’s children based on the requested date. It then selectively explores the respective child nodes that seem relevant to the question using the search function. Note that the LLM is prompted to collapse irrelevant nodes again in order to limit the number of tokens used and speed up further requests. In the given example, when reaching a leaf node, the answer to the question is not evident from any of the natural-language summaries on each level so the LLM decides to invoke a VLM to gather more information. Finally, it invokes the answer function to answer the user’s question.
Our implementation of the LLM agent uses a prompting style inspired by the simulated Python console of [14]. The LLM can issue any command – including compound statements such as loops – using the provided API. After the execution of the respective code, the LLM can “see” the output of its command(s), or any execution error. This process is repeated, and the prompt to the LLM always contains the (growing) execution history. Our API returns detailed error messages and usage hints in case the LLM uses it in a wrong way, and we also detect repetitive generations, prompting the LLM to try to solve the problem differently. In case an answer is not obtained after a set amount of maximum iterations, we use a special prompt to force an immediate answer given the current state of the history tree. Zero-shot experiments prompt the LLM with only a static prefix to explain the task and the available API, while few-shot prompting adds top-$k$ examples selected based on semantic similarity to the current user query. The final part of the prompt is always a string representation of the history tree’s current state.
@@ IV-A Simulated Household Episodes
Data:   Episodes in TEACh are on average $6.2\pm 5.3$ min long. Since we are interested in very long histories of robot experience, we randomly combine them to form histories of up to 100 episodes. We also randomize dates and times for each episode, ensuring realistic sequences by picking one to five episodes per day, avoiding nighttimes, and occasionally skipping some days; the longest histories thus span nearly two months. Based on these histories, we generate QA pairs by adjusting the generation grammar from [7]. Specifically, we generate ten types of questions. These ask for: a list of high-level summaries of episodes (task descriptions); a detailed description of one particular episode in a history; a summary of an episode that happened either before or after a particular episode; a list of episodes in which a particular object was seen or action performed; a summary of an episode that occurred at a given time or a specified number of days ago; a list of times or number of days ago at which a given task was performed. Questions can refer to specific episodes either by date and time, or by the task performed (if it is unique). From the TEACh “valid unseen” set, we generate test sets with 10 histories per sequence length (combining |h|=5, 15, 25, 50|h|=5,\,15,\,25,\,50, and $100$ episodes). Each history $h$ is annotated with ten QA pairs, making up 100 samples per history length $|h|$.
TABLE I: Results on Simulated Household Episodes from TEACh [17]
@@
img 1-pass FS flat 24 49 164 13 43 334 17 41 406 11 35 512 10 30 1256
text 1-pass 0 L3 5 18 65.4 8 15 336 5 15 270 5 25 789 6 14 1234
H-Emv 1 full 9 40 10.6 9 37 10.2 10 42 10.2 15 36 11.6 15 35 11.9
@@
img 1-pass FS flat 57 83 165 30 73 336 39 69 418 25 60 861 17 54 1591
text 1-pass 0 L3 21 43 68.9 19 43 217 22 50 287 15 41 740 7 28 1304
H-Emv 1 full 28 64 9.2 33 63 9.9 31 63 9.7 27 57 10.9 19 51 13.3
@@
ReMEmbR 10 30 12.2 6 21 12.5 11 28 12.5 7 22 12.6 5 18 12.7
text 1-pass 0 L3 32 65 120 37 63 372 35 69 422 24 56 1055 OOC
H-Emv 0 L3 28 61 19.7 20 62 44.2 29 65 19.8 16 64 116 21 52 249
H-Emv 0 full 51 71 3.1 45 76 4.1 46 73 15.0 32 58 4.9 25 60 5.6
H-Emv 1 full 48 74 8.5 48 72 9.6 55 74 23.7 37 68 10.2 34 62 10.4
$S_{c},S_{p}$: semantically categorized (partially) correct in %, T: number of 1K prompt token.
OOC: out of context, ICL: in-context learning, FS: few-shot
$10^{1}$$10^{2}$$10^{3}$0%2020%4040%6060%8080%100100%1K Tokens per sample ($T$)H-Emv 1-shotH-Emv 0-shotH-Emv (L3)text 1-pass L3
Figure 4: Token costs vs. performance for different history lengths (TEACh full multimodal). Solid lines are $S_{p}$, dashed lines $S_{c}$. H-Emv retains better performance at lower costs.
@@
Evaluation Metrics:   Evaluation of free-form EMV answers is hard since there can be many ways to formulate the correct answer, questions can be underspecified, and verifying abstract statements by grounding them in the history tree is a research question in itself. Following [1], we define a semantic categorization of a model’s output $o$ given the GT $g$ and question $q$: correct when $o$ is semantically equivalent to $g$; correctly summarized if $o$ is a correctly summarized version of $g$, still containing all relevant facts (in context of $q$); correct TMI (too much information) if $o$ is correct but overly specific; partially correct TMI if parts of $o$ are correct, but there are TMI parts and these are wrong; partially correct missing if parts of $o$ are correct, but relevant facts from $g$ are missing; wrong when $o$ could be an answer to $q$ but is none of the above; and no answer if $o$ is empty, completely irrelevant to $q$, or the model threw an error. Counting correct (incl. correct summarized, correct TMI) samples $C$ and at least partially correct samples $P$ ($C\subseteq P$), we report their percentages $S_{c}$ and $S_{p}$, respectively. Since categorizing each evaluated sample by hand is prohibitively expensive, we prompt GPT-4o [56] to perform this evaluation. For this, we started by annotating 60 samples by hand, and use these as a database to retrieve few-shot samples based on maximal marginal relevance [57]. We further tuned the prompts on a validation set of 100 hand-categorized model outputs (generated from TEACh train).
@@
We evaluate the agreement of the LLM’s predicted categories with manually annotated ones on 200 model results from our test data, resulting in an aggregated category accuracy of $88\%$, and per-class f-scores of $F_{1}(\text{correct})=0.89,F_{1}(\text{partially\_correct})=0.84,F_{1}(\text{wrong})=0.91$. The LLM categorizes correct and wrong samples very well and has the most difficulties on the partially correct labels. However, these categories are also defined imprecisely, and the inter-annotator agreement [58] between the first two authors has only a value of Cohen’s $\kappa=0.66$ ($n=110$), vs. $\kappa=0.91$ ($n=68$) when only considering correct/wrong. Thus, while not perfect, we use the LLM to automatically obtain reasonable score estimates. We do not report surface-level automated metrics such as BLEU [59] or ROUGE [60] as these metrics struggle to capture the variety of correct answers in the EMV task, e. g., for a “when”-question, both of the following answers are correct, but have no word overlap at all: “at 4 PM” vs. “in the afternoon”.
@@
Settings and Baselines:   We evaluate under three settings: First, vision-only can act solely on the visual data stream. As a baseline, we prompt Gemini 1.5 Pro [11] in one pass with the sequence of images along with timestamps and the question (img 1-pass). We choose Gemini because we require extremely long requests; however, despite its 2M token limit, we need to sample every 2nd frame so that longer histories fit into the context window. We few-shot-prompt with one static example history of five episodes including ten QA samples for this history. For a fair comparison, we use the same LLM for ReMEmbR and H-Emv (although we can achieve better results with GPT-4o, as shown in Sec. IV-C). H-Emv does not take raw images, but constructs history trees by inferring objects using YOLO-World [61] and actions using a LongT5 transformer model [62] fine-tuned on TEACh train. This vision model limits our system with an action classification accuracy of $62\%$ (which can be partially circumvented by the summarization LLM’s commonsense knowledge). Few-shot prompting Gemini to infer the sequence of actions for building the tree had a worse accuracy of around $27\%$. We also compare with one-pass zero-shot prompting the LLM with this tree limited to L3 node (i. e., no hierarchical summarization) in text form (text 1-pass). Second, vision + speech enriches the visual information with the dialog data from TEACh episodes, representing natural language commands given to the robot. This is simply added to the prompt for img 1-pass, and inserted into the history tree for H-Emv and text 1-pass. Finally, full multimodal uses the recorded (GT) actions and goals from the TEACh episodes, as this information is typically available when a robot executes some actions. This setting also uses GT object information to compare system performance assuming perfect vision components. We compare H-Emv with one-shot and zero-shot prompting of the LLM agent. For preparing the few-shot samples, we use histories built from episodes in TEACh train, and record traces of manually using the Python console interface and the defined API to interact with the history tree until the given GT answer becomes evident. While we collect two to three samples per question type this way, making up 21 samples in total, we select only the top-1 sample when prompting the LLM, based on semantic similarity of the user’s questions. Semantic similarity is determined after asking gpt-4o-mini to cross out the task-specific words from the question so that an example from the same question type is retrieved (instead of an irrelevant example just mentioning the same objects or activities). We further ablate the hierarchical summarization: H-Emv 0-shot L3 is our method without LLM-generated summaries (L4+), still using the interactive agent to explore an initially collapsed list of L3 nodes. Last, text 1-pass 0-shot L3 is a baseline presenting the fully expanded tree (L3 and lower) to the LLM along with the question in a single prompt, thus not using the LLM as an agent. We also compare to ReMEmbR [28], with their method/prompting minimally adjusted to our setting.
Results:   Results of our TEACh experiments can be found in Fig. 4. First, we can observe that every method’s performance decreases with increasing $|h|$.
@@
Focusing on the vision-only and vision + speech results, the Gemini 1-pass baseline outperforms H-Emv for shorter histories. This is reasonable, as the baseline can directly access the full stream of visual information, whereas our hierarchical system suffers from error propagation and is limited by pretrained vision components. In particular, the history tree could contain incomplete or wrong information or our method could fail by expanding the wrong nodes of the tree, which cannot happen to the 1-pass baseline. However, token costs scale linearly with history length for 1-pass, while it stays approximately constant for H-Emv. The performance also drops faster for 1-pass, with H-Emv reaching comparable or better semantic scores for $|h|\geq 25$. In contrast, when circumventing the limitations of perception components by using GT object detection and action information (full multimodal setting), H-Emv outperforms the 1-pass system with a token budget two orders of magnitudes smaller (see Fig. 4). Further, 1-shot prompting helps for longer histories, but 0-shot works with half the token costs and comparable performance for short histories. Ablating the hierarchical higher-level summaries notably increases token cost and leads to worse performance, especially for longer histories (when the LLM sometimes expands all nodes). While ReMEmbR keeps token cost low due to its use of the LLM as an agent similar to H-Emv, its retrieval-augmented generation approach and lack of summarization capability has difficulties with our QA data, as it typically retrieves single actions (“I picked up a mug”) instead of higher-level task information (“I prepared breakfast”).